# First Statistical Detection of Cool Gas Outflows with JWST Towards Cosmic Dawn

### v1

@@ Abstract
Galactic-scale outflows are a crucial component of galaxy evolution, yet their properties in the early universe remain poorly constrained. We present the first statistical investigation of cool gas outflows in galaxies spanning a wide cosmic timeline from $z\approx 1$ to $z\approx 10$. Using thousands of public JWST/NIRSpec spectra, we employ a signal-to-noise weighted spectral stacking technique on the Mg II $\lambda\lambda 2796,2803$ absorption doublet. We robustly detect blueshifted Mg II absorption in all stellar mass and redshift bins. The outflow equivalent width exhibits a strong, positive correlation with stellar mass ($M_{*}$) at all epochs, increasing from $\sim 1$ Å at $M_{*}\approx 10^{9}~\mathrm{M}_{\odot}$ to over $3$ Å at $M_{*}>10^{10.5}~\mathrm{M}_{\odot}$. Our work provides the first statistical constraints on cool outflows in the low-mass ($M_{*}\lesssim 10^{9.5}~\mathrm{M}_{\odot}$), high-redshift ($z>3$) regime, vital for constraining feedback in the numerous progenitors of typical present-day galaxies. Crucially, the scaling relation between outflow properties and stellar mass shows no significant evolution at $z>3$. This suggests a persistent, unevolving feedback mechanism governing the baryon cycle in the early universe, placing strong constraints on models that invoke a fundamental change in feedback physics at Cosmic Dawn, such as the feedback-free starburst model.
@@ 1 Introduction
Galactic-scale outflows are fundamental to our understanding of galaxy evolution. Driven by energetic feedback from massive stars, supernovae (SNe), and active galactic nuclei (AGN) (e.g., Silk1998; Veilleux2005; Croton2006; Veilleux2020), these outflows act as a key regulatory agent in the baryon cycle between galaxies and their surrounding environments. By expelling gas from the interstellar medium (ISM), they modulate the fuel available for star formation, thereby shaping the mass assembly of galaxies over cosmic time (e.g., DiMatteo2005; Hopkins2008; Somerville2008; Erb2015). Furthermore, these outflows are the primary mechanism for transporting chemically enriched material from galaxies into the circumgalactic medium (CGM) and even the intergalactic medium (IGM), playing a crucial role in the cosmic metal enrichment history (e.g., Pettini2001; Tremonti2004; Finlator2008; Tumlinson2017; Wang2021; Chen2025a; Chen2025b; Lyu2025a). Despite their acknowledged importance in theoretical models and hydrodynamical simulations of galaxy formation, establishing robust observational constraints on the physical properties of outflows remains a significant challenge, particularly in the early universe.
@@
Observing outflows becomes progressively more difficult at high redshifts. During the peak epoch of cosmic star formation ($z\sim 1-3$), outflows are known to be ubiquitous in massive star-forming galaxies (e.g., Weiner2009; Rubin2014; ForsterSchreiber2019). However, pushing these studies towards the epoch of Cosmic Dawn ($z>6$) presents formidable challenges. The intrinsic faintness of high-redshift galaxies makes it extremely difficult to obtain spectra with sufficient signal-to-noise ratio (S/N) to detect the subtle absorption features imprinted by outflows on the galaxy’s continuum. Consequently, most studies have been limited to either lensed systems or the most luminous, rare objects (e.g. Valentino2025), hindering a comprehensive statistical analysis.
Absorption line spectroscopy is a powerful and widely used technique to probe the kinematics of cool gas flows along the line of sight. Blueshifted low-ionization metal absorption lines, such as the Mg II $\lambda\lambda 2796,2803$ doublet and the Na I D $\lambda\lambda 5890,5896$ doublet, are particularly effective tracers of cool ($T\sim 10^{4}$ K), neutral gas in outflows (e.g., Weiner2009; Martin2012; Rubin2014). The Mg II doublet, commonly detected in the CGM of galaxies, is a robust indicator of enriched cool gas over a wide range of H I column densities (Bergeron1986; Churchill2000). The Na I D doublet, with its low ionization potential (5.1 eV), specifically traces dense, neutral regions often shielded by dust (Savage1996; Baron2020). However, due to their weakness, detecting these features in individual high-redshift galaxies is often infeasible.
The advent of the James Webb Space Telescope (JWST, Gardner2006) is revolutionizing the study of high-redshift galaxies. With its unprecedented sensitivity and near-infrared spectroscopic capabilities, JWST/NIRSpec provides high-quality spectra that are essential for outflow studies, free from the bright atmospheric sky background that hampers ground-based observations. This enables the detection of rest-frame optical features like Na I D at high redshifts for the first time (e.g., Belli2024; Cresci2023; Davies2024; Kehoe2025), and allows for the use of consistent rest-frame UV tracers like Mg II across a vast cosmic timeline ($z\sim 1-10$). The ability to assemble large spectroscopic samples of typical galaxies at these early epochs finally opens the door to characterizing the statistical properties of outflows and establishing crucial scaling relations that link outflow properties to intrinsic galaxy parameters, such as stellar mass and redshift.
JWST also opens a new frontier for testing galaxy formation theories in the early universe. Recently, new models have been proposed to explain the surprising abundance of luminous galaxies observed by JWST at $z\gtrsim 10$. One prominent example is the “feedback-free starburst” (FFB) model (Dekel2023; Li2024). This model posits that at the high gas densities and low metallicities prevalent at Cosmic Dawn, the star formation timescale can become shorter than the delay time for stellar feedback from supernovae and stellar winds. This could lead to a highly efficient, bursty mode of star formation, fundamentally different from the self-regulating feedback processes observed at later cosmic times. The physical properties of galactic outflows—such as their velocities, equivalent widths, and scaling relations with galaxy properties—serve as direct probes of the underlying feedback mechanisms. By statistically characterizing outflows towards Cosmic Dawn, we can test the predictions of models like the FFB and determine whether a shift in the nature of feedback indeed occurred in the early universe.
In this work, we leverage the vast public dataset from the DAWN JWST Archive (DJA) to conduct the statistical investigation of cool gas outflows in galaxies spanning the redshift range $1<z<10$. By employing a spectral stacking technique on hundreds of JWST/NIRSpec spectra, we significantly enhance the S/N, enabling the detection of the faint, average outflow signals that are invisible in individual spectra. The primary goals of this Letter are: (1) to reliably detect outflow traced by Mg II and Na I D at different redshift, and measure the average outflow velocities and equivalent widths (EW) traced by Mg II across a range of redshifts and stellar masses; and (2) to quantitatively establish the scaling relations between outflow properties, galaxy stellar mass, and cosmic time; and (3) to provide key observational constraints on the physics of galactic feedback and its role in galaxy evolution towards the Cosmic Dawn.
@@
2 Data and Sample Selection
@@
Figure 1: Sample selection and stellar mass distribution. The left panel displays the distribution of our galaxy samples in the stellar mass-redshift plane. The grey contours represent the parent sample of 24,927 galaxies, while the blue and red contours show the subsamples with valid spectral coverage of the Mg II and Na I D doublets, respectively. The three contour levels for each sample are derived from a Kernel Density Estimate (KDE) and correspond to the 16th, 50th, and 84th percentiles of the distribution. The middle and right panels present the Probability Density Functions (PDFs) of the stellar mass distributions for the Mg II- and Na I D-available subsamples. The colored histograms correspond to the different redshift bins as labeled. To assess the representativeness of our sample, we overlay the evolving Galaxy Stellar Mass Function (GSMF) from Weaver2023. The brown shaded region represents the envelope of the GSMF across the redshift range $1.1<z<7.5$, derived from the Schechter function parameters (double for $z<3$ and single for $z>3$) presented in their work. The close agreement in shape demonstrates that our sample is broadly representative of the underlying galaxy population at these epochs.
@@
Our analysis is based on the JWST/NIRSpec spectra dataset from the DAWN JWST Archive (DJA) [^1] https://dawn-cph.github.io/dja/. Specifically, we utilize the v4.4 Merged Table  [^2] https://dawn-cph.github.io/dja/blog/2025/05/01/nirspec-merged-table-v4/ and its associated extended spectra catalog (Valentino2025; Pollock2025). This comprehensive public catalog provides 80,367 uniformly processed spectra, reduced using the msaexp (Brammer2022; deGraaff2025; Heintz2025) and grizli (Brammer2023) pipelines. The data products include 1D/2D spectral extractions, redshift measurements, emission-line fluxes, and associated photometry. The complete dataset is available at DOI: 10.5281/zenodo.1547235.
@@
We construct our parent sample by selecting galaxies with a photometrically derived stellar mass ($M_{*}$) greater than $10^{8}~M_{\odot}$ and a secure spectroscopic redshift. The redshift quality is assessed via visual inspection flags provided in the catalog. We exclude spectra with poor quality flags, specifically Grade=0=0 (spectrum suffers some data quality issue) and Grade=1-1 (fit not performed or graded). This initial selection results in a sample of 24,927 galaxy spectra.
@@
To search for cool gas outflows, we create subsamples based on the spectral coverage of the Mg II $\lambda\lambda 2796,2803$ and Na I D $\lambda\lambda 5890,5896$ doublets. We select objects observed with NIRSpec medium- or high-resolution grisms (G140M/H, G235M/H, G395M/H; $R\sim 1000/2700$) that provide unmasked spectral coverage of a 100 Å window roughly centered on these doublets (i.e., 2750–2850 Å for Mg II and 5850–5950 Å for Na I D). This selection ensures that the absorption features, if present, fall within the observed spectral range and are sampled at a sufficient resolution for detection. This procedure yields final samples of approximately 2,394 spectra for the Mg II analysis and 5,512 spectra for the Na I D analysis. However, due to the emission line blending and inherent stellar contamination, we do not investigate the scaling relations with galaxy properties of Na I D sample in this work (See Section 4.1 for details).
@@
To investigate whether our sample is a representative galaxy sample, we compare our distributions to the evolving Galaxy Stellar Mass Function (GSMF) from the comprehensive JWST analysis of Weaver2023. Although our sample is from a wide range of programs and surveys for different propose, as shown by the brown shaded envelope in Figure 1, which encapsulates the GSMF evolution from $z=1.1$ to $z=7.5$, our sample distributions trace the shape of the underlying galaxy population remarkably well, particularly at the high-mass end ($M_{*}>10^{9}\rm{M_{\odot}}$) where the majority of the cosmic stellar mass resides. This agreement confirms that our spectroscopic sample is broadly representative of the typical galaxy population across the full mass and redshift range of our study. Therefore, the outflow properties derived in this work are characteristic of the general galaxy population during these cosmic epochs.
3 Spectra Stacking Method
To systematically investigate the prevalence of galactic outflows across the galaxy population, we stack galaxy spectra binned by their physical properties. This technique significantly improves S/N, enabling the detection of outflow signatures that are too faint to identify in individual spectra (e.g. Weiner2009; Chen2010; Rubin2010; Bordoloi2014).
3.1 Weighted Median Stacking
We begin by shifting all individual spectra to their rest-frame. For each spectrum, we extract the spectral data within a 200Å window centered on the doublet features (2700–2900Å for Mg II and (5800–6000Å for Na I D) . These segments are then resampled onto a unified wavelength grid using Spectres (Carnall2017). To avoid artificial super-sampling, which could introduce additional errors, the wavelength interval for the resampled spectra is inferred from the sample redshift, resulting in adopted intervals of $\Delta\lambda=1.4\ {\rm\AA },0.9\ {\rm\AA },0.7\ {\rm\AA },0.6\ {\rm\AA },0.5\ {\rm\AA }$ for $1<z<2$, $2<z<3$, $3<z<4$, $4<z<5$ and $5<z<10$ bins, respectively. After masking the prominent lines (Mg II, Mg I, Na I D), we estimate the continuum by first applying a median filter with an 11-pixel window and then fitting a 5th-order polynomial to the smoothed spectrum. Each original spectrum is divided by its corresponding continuum to produce a normalized spectrum. In idealized circumstances, the normalized flux should be unity across non-absorbing regions. However, the normalization can be biased in cases of low S/N or unexpected strong absorption features, thereby introducing additional uncertainty into the stacked spectrum. For instance, if the fitted continuum approaches or crosses zero, the normalized flux can diverge, producing a singularity.
@@ 3.1 Weighted Median Stacking
where $F_{i}$ is the $i$th flux density in the normalized spectrum, $F_{i}^{\rm ref}=1$ is the reference value for a perfectly normalized spectrum, and $N$ is the quantity of the spectral data points. In this calculation, the regions with expected spectral lines (Mg II, Mg I, Na I D) are masked. We then generate the median-stacked spectra based on these weights. The composite flux in each wavelength bin is computed as the weighted median of all individual fluxes. For $n$ ordered fluxes $f_{1},f_{2},\cdots,f_{n}$ with corresponding weights $w_{1},w_{2},\cdots,w_{n}$, the weighted median is defined as the element $f_{k}$ satisfying
@@
To quantify the uncertainty in the final stacked spectra, we employ a bootstrap resampling procedure. Specifically, we generate 1000 bootstrap samples, each by randomly drawing $N$ spectra from the original set of $N$ with replacement. For comparison, we have also performed an unweighted stacking analysis with equal probabilities; the results are presented in the Appendix A.
@@
3.2 Outflow Properties
We employ the boxcar method (Rubin2010; Bordoloi2014) to derive the mean outflow velocity ($v_{\rm out}$) and the outflow equivalent width ($\rm EW_{out}$) of the Mg II doublet. This approach provides model-independent measurements directly from the observed absorption profile, which is advantageous for our statistical analysis. Given the blending between the red-wing of Mg II $\lambda$2796 line and the blueshifted Mg II $\lambda$2803 absorption, we use the blue-wing of Mg II $\lambda$2796 and the red-wing of Mg II $\lambda$2803 to trace the kinematics of the species. We conceptually decompose the absorption into a blueshifted outflowing component (denoted as “out”) characterized by a blueshifted line center, and an interstellar component (denoted as “sym”) with a symmetrical profile centered at the rest-frame wavelength. The EWs are measured in two velocity intervals: $\rm EW_{blue}$ from 700-700 to $0~\rm km\ s^{-1}$ relative to Mg II $\lambda$2796, and $\rm EW_{red}$ from 0 to $400~\rm km\ s^{-1}$ relative to Mg II $\lambda$2803, which are defined empirically to cover the extended absorption wings. Following Bordoloi2014, the total absorption EW, defined as $\rm EW_{tot}\equiv EW_{out}+EW_{sym}$, and the outflow EW are estimated as

### v2

# First Statistical Detection of Mg II-traced Cool Gas Outflows with JWST Towards Cosmic Dawn

@@ Abstract
Galactic-scale outflows are a crucial component of galaxy evolution, yet their properties in the early universe remain poorly constrained. We present the first statistical investigation of Mg II-traced cool gas outflows in galaxies spanning a wide cosmic timeline from $z\approx 1$ to $z>6$ (with sample coverage extending to $z\sim 10$). Using thousands of public JWST/NIRSpec spectra, we employ a signal-to-noise weighted spectral stacking technique on the Mg II $\lambda\lambda 2796,2803$ absorption doublet. We robustly detect blueshifted Mg II absorption in nearly all stellar mass and redshift bins, with the exception of the lowest-mass systems at $z\sim 1-2$. The outflow equivalent width exhibits a positive correlation with stellar mass ($M_{*}$) at all epochs, with the fitted slope of $1.21\pm 0.35$. Our work provides the first statistical constraints on Mg II-traced cool outflows in the low-mass ($M_{*}\lesssim 10^{9.5}\,\mathrm{M}_{\odot}$), high-redshift ($z>3$) regime. We also find that the outflow velocities generally remain below the host halo escape velocities, consistent with a galactic fountain scenario. The consistency of the stellar mass-outflow equivalent width relation across $z\sim 2-6$ suggests a persistent, unevolving feedback mechanism governing the baryon cycle towards cosmic dawn.
@@ I Introduction
Galactic-scale outflows are fundamental to our understanding of galaxy evolution. Driven by energetic feedback from massive stars, supernovae (SNe), and active galactic nuclei (AGN) (e.g., Silk and Rees, 1998; Veilleux et al., 2005; Croton et al., 2006; Veilleux et al., 2020), these outflows act as a key regulatory agent in the baryon cycle between galaxies and their surrounding environments. By expelling gas from the interstellar medium (ISM), they modulate the fuel available for star formation, thereby shaping the mass assembly of galaxies over cosmic time (e.g., Di Matteo et al., 2005; Hopkins et al., 2008; Somerville et al., 2008; Erb, 2015; Wang et al., 2019). Furthermore, these outflows are the primary mechanism for transporting chemically enriched material from galaxies into the circumgalactic medium (CGM) and even the intergalactic medium (IGM), playing a crucial role in the cosmic metal enrichment history (e.g., Pettini et al., 2001; Tremonti et al., 2004; Finlator and Davé, 2008; Tumlinson et al., 2017; Wang and Lilly, 2021, 2022; Ma et al., 2024; Chen et al., 2025b, a; Jia et al., 2025; Lyu et al., 2025b). Despite their acknowledged importance in theoretical models and hydrodynamical simulations of galaxy formation, establishing robust observational constraints on the physical properties of outflows remains a significant challenge, particularly in the early universe.
@@
Observing outflows becomes progressively more difficult at high redshifts. During the peak epoch of cosmic star formation ($z\sim 1-3$), outflows are known to be ubiquitous in massive star-forming galaxies (e.g., Weiner et al., 2009; Rubin et al., 2014; Förster Schreiber et al., 2019). However, pushing these studies towards the epoch of cosmic dawn ($z>6$) presents formidable challenges. The intrinsic faintness of high-redshift galaxies makes it extremely difficult to obtain spectra with sufficient signal-to-noise ratio (S/N) to detect the subtle absorption features imprinted by outflows on the galaxy’s continuum.
@@
Absorption line spectroscopy is also a powerful and widely used technique to probe the kinematics of cool gas flows along the line of sight. Blueshifted low-ionization metal absorption lines, such as the Mg II $\lambda\lambda 2796,2803$ doublet and the Na I D $\lambda\lambda 5890,5896$ doublet, are particularly effective tracers of cool, neutral gas in outflows (e.g., Weiner et al., 2009; Martin et al., 2012; Rubin et al., 2014). The Mg II doublet, commonly detected in the CGM of galaxies, is a robust indicator of enriched cool gas ($T\sim 10^{4}$ K) over a wide range of H I column densities (Bergeron and Stasińska, 1986; Churchill et al., 2000). The Na I D doublet, with its low ionization potential (5.1 eV), specifically traces dense, neutral gas ($T<10^{3}$ K) often shielded by dust (Savage and Sembach, 1996; Puspitarini and Lallement, 2012; Baron et al., 2020). However, detecting these subtle absorption features in individual high-redshift galaxies remains challenging. Valentino et al. (2025) recently reported individual detections of blueshifted Mg II absorption in two rare massive quiescent galaxies at $z\sim 4$ and $z\sim 7$. However, a statistical census of the cool, low-ionization gas phase traced by Mg II in the typical, abundant star-forming galaxy population remains largely unexplored.
JWST/NIRSpec enables the detection of rest-frame optical features like Na I D at high redshifts for the first time (e.g., Belli et al., 2024; Cresci et al., 2023; Davies et al., 2024; Kehoe et al., 2025), and UV tracers like Mg II across a vast cosmic timeline ($z\sim 1-10$). The ability to assemble large spectroscopic samples of typical galaxies at these early epochs finally opens the door to characterizing the statistical properties of outflows and establishing crucial scaling relations that link outflow properties to intrinsic galaxy parameters, such as stellar mass and redshift.
JWST also opens a new frontier for testing galaxy formation theories in the early universe. Recently, new models have been proposed to explain the surprising abundance of luminous galaxies observed by JWST at $z\gtrsim 10$. One prominent example is the “feedback-free starburst” (FFB) model (Dekel et al., 2023; Li et al., 2024). This model posits that at the high gas densities and low metallicities prevalent at cosmic dawn, the star formation timescale can become shorter than the delay time for stellar feedback from supernovae and stellar winds. This could lead to a highly efficient, bursty mode of star formation, fundamentally different from the self-regulating feedback processes observed at later cosmic times. The physical properties of galactic outflows—such as their velocities, equivalent widths, and scaling relations with galaxy properties—serve as direct probes of the underlying feedback mechanisms. By statistically characterizing outflows towards cosmic dawn, we can test the predictions of models like the FFB and determine whether a shift in the nature of feedback indeed occurred in the early universe.
In this work, we leverage the vast public dataset from the DAWN JWST Archive (DJA) to conduct the statistical investigation of cool gas outflows in galaxies spanning the redshift range $1<z<10$. Note that while our sample includes galaxies out to $z\approx 10$, the S/N-weighted stacking in the highest redshift bin is most sensitive to the population at $z\sim 6$, see Table 1 for weighted mean redshifts. By employing a spectral stacking technique on hundreds of JWST/NIRSpec spectra, we significantly enhance the S/N, enabling the detection of the faint, average outflow signals that are invisible in individual spectra. The primary goals of this Letter are: (1) to reliably detect outflow traced by Mg II and Na I D at different redshift, and measure the average outflow velocities and equivalent widths traced by Mg II across a range of redshifts and stellar masses; and (2) to quantitatively establish the scaling relations between outflow properties, galaxy stellar mass, and across cosmic time; and (3) to provide key observational constraints on the physics of galactic feedback and its role in galaxy evolution towards the cosmic dawn.
[…]
II Data and Sample Selection
@@
Figure 1: Sample selection and stellar mass distribution. The left panel displays the distribution of our galaxy samples in the stellar mass-redshift plane. The grey contours and scatter points represent the parent sample of 24,927 galaxies, while the blue and red contours and points show the subsamples with valid spectral coverage of the Mg II and Na I D doublets, respectively. The three contour levels for each sample are derived from a Kernel Density Estimate (KDE) and correspond to iso-density boundaries enclosing 16%, 50%, and 84% of the distribution. Individual galaxies located in low-density regions (outside the 84% contour) are plotted as scatter points to illustrate the full extent of the sample. The middle and right panels present the Probability Density Functions (PDFs) of the stellar mass distributions for the Mg II- and Na I D-available subsamples. The colored histograms correspond to the different redshift bins as labeled. To assess the representativeness of our sample, we overlay the evolving Galaxy Stellar Mass Function (GSMF) from Weaver et al. (2023). The brown shaded region represents the envelope of the GSMF across the redshift range $1.1<z<7.5$, derived from the Schechter function parameters (double for $z<3$ and single for $z>3$) presented in their work.
@@
Our analysis is based on the JWST/NIRSpec spectra dataset from the DAWN JWST Archive (DJA) [^1] https://dawn-cph.github.io/dja/. Specifically, we utilize the v4.4 Merged Table  [^2] https://dawn-cph.github.io/dja/blog/2025/05/01/nirspec-merged-table-v4/ and its associated extended spectra catalog (Valentino et al., 2025; Pollock et al., 2025). This comprehensive public catalog provides 80,367 uniformly processed spectra, reduced using the msaexp (Brammer, 2022; de Graaff et al., 2025; Heintz et al., 2025) and grizli (Brammer, 2023) pipelines. The data products include 1D/2D spectral extractions, redshift measurements, emission-line fluxes, and associated photometry. The complete dataset is available at DOI: 10.5281/zenodo.1547235.
@@
We construct our parent sample by selecting galaxies with a photometrically derived stellar mass ($M_{*}$) greater than $10^{8}~M_{\odot}$ and a secure spectroscopic redshift. The redshift quality is assessed via visual inspection flags provided in the catalog. We exclude spectra with poor quality flags, specifically Grade=0=0 (spectrum suffers some data quality issue) and Grade=1-1 (fit not performed or graded). We note that the vast majority (approximately 99.5%) of our selected galaxies are assigned Grade 3, indicating robust redshifts derived from multiple high-S/N emission lines. For these sources, the typical redshift uncertainty corresponds to a velocity error of $\lesssim 30$ km s-1, which is negligible compared to the typical outflow velocities derived in this work. This initial selection results in a sample of 24,927 galaxy spectra.
@@
To search for cool gas outflows, we create subsamples based on the spectral coverage of the Mg II $\lambda\lambda 2796,2803$ and Na I D $\lambda\lambda 5890,5896$ doublets. We select objects observed with NIRSpec medium- or high-resolution grisms (G140M/H, G235M/H, G395M/H; $R\sim 1000/2700$) that provide unmasked spectral coverage of a 100 Å window roughly centered on these doublets (i.e., 2750–2850 Å for Mg II and 5850–5950 Å for Na I D). This selection ensures that the absorption features, if present, fall within the observed spectral range and are sampled at a sufficient resolution for detection. This procedure yields final samples of approximately 2,394 spectra for the Mg II analysis and 5,512 spectra for the Na I D analysis. However, due to the emission line blending and inherent stellar contamination, we do not investigate the scaling relations with galaxy properties of Na I D sample in this work (See Section IV.1 for details).
@@
To investigate whether our sample is a representative galaxy sample, we compare our distributions to the evolving Galaxy Stellar Mass Function (GSMF) from the comprehensive JWST analysis of Weaver et al. (2023). Although our sample is from a wide range of programs and surveys for different proposes, as shown by the brown shaded envelope in Figure 1, which encapsulates the GSMF evolution from $z=1.1$ to $z=7.5$, our sample distributions trace the shape of the underlying galaxy population remarkably well, particularly at the high-mass end ($M_{*}>10^{9}\rm{M_{\odot}}$) where the majority of the cosmic stellar mass resides. This agreement confirms that our spectroscopic sample is broadly representative of the typical galaxy population across the full mass and redshift range of our study. Therefore, the outflow properties derived in this work are characteristic of the general galaxy population during these cosmic epochs.
III Spectra Stacking Method
To systematically investigate the prevalence of galactic outflows across the galaxy population, we stack galaxy spectra binned by their physical properties. This technique significantly improves S/N, enabling the detection of outflow signatures that are too faint to identify in individual spectra (e.g. Weiner et al., 2009; Chen et al., 2010; Rubin et al., 2010; Bordoloi et al., 2014).
III.1 Weighted Median Stacking
We begin by shifting all individual spectra to their rest-frame. For each spectrum, we extract the spectral data within a 200Å window centered on the doublet features (2700–2900Å for Mg II and 5800–6000Å for Na I D). These segments are then resampled onto a unified wavelength grid using Spectres (Carnall, 2017). To avoid artificial super-sampling, which could introduce additional errors, the wavelength interval for the resampled spectra is inferred from the sample redshift, resulting in adopted intervals of $\Delta\lambda=1.4\ {\rm\AA },0.9\ {\rm\AA },0.7\ {\rm\AA },0.6\ {\rm\AA },0.5\ {\rm\AA }$ for $1<z<2$, $2<z<3$, $3<z<4$, $4<z<5$ and $5<z<10$ bins, respectively. After masking the prominent lines (Mg II, Mg I, Na I D), we estimate the continuum by first applying a median filter with an 11-pixel window and then fitting a 5th-order polynomial to the smoothed spectrum. Each original spectrum is divided by its corresponding continuum to produce a normalized spectrum. In idealized circumstances, the normalized flux should be unity across non-absorbing regions. However, the normalization can be biased in cases of low S/N or unexpected strong absorption features, thereby introducing additional uncertainty into the stacked spectrum. For instance, if the fitted continuum approaches or crosses zero, the normalized flux can diverge, producing a singularity.
@@ III.1 Weighted Median Stacking
where $F_{i}$ is the $i$th flux density in the normalized spectrum, $F_{i}^{\rm ref}=1$ is the reference value for a perfectly normalized spectrum, and $N$ is the number of the spectral data points. In this calculation, the regions with expected spectral lines (Mg II, Mg I, Na I D) are masked. We then generate the median-stacked spectra based on these weights. The composite flux in each wavelength bin is computed as the weighted median of all individual fluxes. For $n$ ordered fluxes $f_{1},f_{2},\cdots,f_{n}$ with corresponding weights $w_{1},w_{2},\cdots,w_{n}$, the weighted median is defined as the element $f_{k}$ satisfying
@@
To quantify the uncertainty in the final stacked spectra, we employ a bootstrap resampling procedure. Specifically, we generate 1000 bootstrap samples, each by randomly drawing $N$ spectra from the original set of $N$ with replacement.
For comparison, we have also performed an unweighted stacking analysis with equal probabilities; the results are presented in the Appendix A.
@@
III.2 Outflow Properties
We employ the boxcar method (Rubin et al., 2010; Bordoloi et al., 2014) to derive the mean outflow velocity ($v_{\rm out}$) and the outflow equivalent width ($\rm EW_{out}$) of the Mg II doublet. This approach provides model-independent measurements directly from the observed absorption profile, which is advantageous for our statistical analysis. Given the blending between the red-wing of Mg II $\lambda$2796 line and the blueshifted Mg II $\lambda$2803 absorption, we use the blue-wing of Mg II $\lambda$2796 and the red-wing of Mg II $\lambda$2803 to trace the kinematics of the species. We conceptually decompose the absorption into a blueshifted outflowing component (denoted as “out”) characterized by a blueshifted line center, and an interstellar component (denoted as “sym”) with a symmetrical profile centered at the rest-frame wavelength. The EWs are measured in two velocity intervals: $\rm EW_{blue}$ from 700-700 to $0~\rm km\ s^{-1}$ relative to Mg II $\lambda$2796, and $\rm EW_{red}$ from 0 to $400~\rm km\ s^{-1}$ relative to Mg II $\lambda$2803, which are defined empirically to cover the extended absorption wings. Following Bordoloi et al. (2014), the total absorption EW, defined as $\rm EW_{tot}\equiv EW_{out}+EW_{sym}$, and the outflow EW are estimated as